Kilauea in Hawaii is the world’s most continuously active volcano. When we say very active volcanoes, they normally eject red-hot rocks and lava instead of smoke and ash. Suppose at an angle 35.0⁰ above the horizontal; a large rock is ejected from the volcano with a speed of 25 𝑚/𝑠. The rock hits the side of the volcano at an altitude of 20.0 𝑚 lower than its starting point. (a) Calculate the time it takes them to follow its path. (b) What are the magnitude and the direction of the rock’s velocity at impact?


1
Expert's answer
2021-09-27T09:03:13-0400

(a) Find the time to reach the highest point:


"t_u=\frac{v\sin35°}{g}=1.46\text{ s}."

The time to fall from the highest point:


"t_d=\sqrt{\frac{2(d+v^2\sin^2\theta/(2g))}{g}}=2.49\text{ s}."

The total time:


"T=t_u+t_d=3.95\text{ s}."

(b) The horizontal and vertical components of velocity will help us to find the resultant velocity at the impact:


"v_x=v\cos\theta=20.5\text{ m/s}.\\
v_y=gt_d=24.4\text{ m/s}.\\
v=\sqrt{v_x^2+v_y^2}=31.9\text{ m/s}."